OpenStudy (anonymous):

x+2y=0 x+2(-2)=0 please solve

OpenStudy (anonymous):

x=4 y=-2

OpenStudy (anonymous):

from the second equation, rewrite it, and you get: x = 4. plug that into the first equation. 4 + 2y = 0 solve for y: 4 = -2y, you get y = -2. final answers: x = 4, y = -2.

OpenStudy (anonymous):

how about -6+2y=0 0+2y=0 6+2y=0 please solve :) thank you so much

OpenStudy (anonymous):

u will get 3 ans for y

OpenStudy (nikita2):

ask your (other) question on left

OpenStudy (anonymous):

3,0 and -3